zarkam21 (zarkam21):

Use the substitution method to solve the system of equations below. Enter your answer as an ordered pair. y=5x-17 y=x+3

OpenStudy (anonymous):

This is very easy! :) You already have y isolated in each equation, so you can set them equal to each other: 5x-17 = x+3 Now solve that for x.

zarkam21 (zarkam21):

isn;t it 4x=-14

zarkam21 (zarkam21):

isn;t it 4x=-14

zarkam21 (zarkam21):

isn;t it 4x=-14

zarkam21 (zarkam21):

isn;t it 4x=-14

zarkam21 (zarkam21):

isn;t it 4x=-14

zarkam21 (zarkam21):

isn;t it 4x=-14

OpenStudy (anonymous):

Close! The left side is right, check your addition on the right side.

zarkam21 (zarkam21):

isn't it -17+3

zarkam21 (zarkam21):

isn't it -17+3

zarkam21 (zarkam21):

isn't it -17+3

OpenStudy (anonymous):

You need to eliminate -17 on the left side, so to do that, you'll add 17 to both sides: 5x - 17=x + 3 (5x - x) - 17 = (x - x) + 3 4x - 17 = 3 4x - 17 + 17 = 3 + 17 4x = 20 Does that make sense so far?

zarkam21 (zarkam21):

Yes So would x= 5

OpenStudy (anonymous):

Yes! Then you plug that in to either of your equations. Using y = x + 3 would probably be easier.

zarkam21 (zarkam21):

So (5,8) right?=)

OpenStudy (anonymous):

Correct :)